A research group conducted a survey about church attendance. The survey respondents were asked about their church attendance and asked to indicate their age.

Church
Attendance
Age
20 to 29 30 to 39 40 to 49 50 to 59
Yes 32 64 95 73
No 68 86 105 77

Use the sample data to determine whether church attendance is independent of age. Use a 0.05 level of significance.

A.State the null and alternative hypotheses.

H0: Church attendance is independent of age.
Ha: Church attendance is not independent of age. H0: Church attendance is not independent of age.

Ha: Church attendance is independent of age.     H0: Church attendance is not mutually exclusive from age.
Ha: Church attendance is mutually exclusive from age. H0: Church attendance is mutually exclusive from age.
Ha: Church attendance is not mutually exclusive from age.

B. Find the value of the test statistic. (Round your answer to three decimal places.)

C. Find the p-value. (Round your answer to four decimal places.)

p-value =

D. State your conclusion.

Do not reject H0. We conclude that church attendance is not independent of age. Reject H0. We cannot conclude that church attendance is not independent of age.    

Do not reject H0. We cannot conclude that church attendance is not independent of age. Reject H0. We conclude that church attendance is not independent of age.

What conclusion can you draw about church attendance as individuals grow older?

Church attendance decreases as individuals grow older. Church attendance increases as individuals grow older.     Church attendance does not change as individuals grow older. No conclusion can be drawn about church attendance as individuals grow older.
Answer #1

A)
H0: Church attendance is independent of age.
Ha: Church attendance is not independent of age.

B)

Observed Frequencies
age
Church
Attendance
20 to 29 30 to 39 40 to 49 50 to 59 Total
yes 32 64 95 73 264
no 68 86 105 77 336
Total 100 150 200 150 600
Expected Frequencies
age
Church
Attendance
20 to 29 30 to 39 40 to 49 50 to 59 Total
yes 44 66 88 66 264
no 56 84 112 84 336
Total 100 150 200 150 600
(fo-fe)^2/fe
3.272727 0.060606 0.556818 0.742424
2.571429 0.047619 0.4375 0.583333

Test statistic, Chi square = sum(Oi-Ei)^2/Ei = 8.272

C)

df = (r-1)*(c-1) = (2-1)*(4-1) = 3

p-value = 0.041 = CHISQ.DIST.RT(8.27,3)

D)

With chisq(3) = 8.27, p<5%, i reject ho at 5 % level of significance and conclude that Church attendance is not independent of age.

Reject H0. We conclude that church attendance is not independent of age.

E)

looking at the table of observed frequencies. it is observed that:

age
20 to 29 30 to 39 40 to 49 50 to 59
Church Attendance - yes 32 64 95 73

Church attendance increases as individuals grow older.